OpenStudy (anonymous):

Y+4=-4(x-2) what point does this line pass through, which is the basis of this equation.

OpenStudy (amistre64):

you have a point slope form of a line

OpenStudy (amistre64):

the basis point is used in the equation itself as the add ons to x and y

OpenStudy (amistre64):

negate the addons, and you have the point

OpenStudy (anonymous):

So the answer would be (4,2)

OpenStudy (amistre64):

almost .... but not quite

OpenStudy (amistre64):

since a point is defined as (x,y) lets start with the x part, what is added onto the x part inthe equation?

OpenStudy (amistre64):

y+4=-4(x-2) ^^ this part right here

OpenStudy (anonymous):

2 or x or -4

OpenStudy (amistre64):

im not sure that that response means

OpenStudy (anonymous):

I'm not sure wat the was so ii put down all possiblities.

OpenStudy (amistre64):

:) ok the portion that is added onto the x part y+4=-4(x-2) ^^ is this part right here, -2 the negation of -2 is: 2 the portion that is added onto the y part y+4=-4(x-2) ^^ is this part right here, +4 the negation of +4 is: -4 can you reconstruct the basis point now?

OpenStudy (anonymous):

2,-4

OpenStudy (amistre64):

that would be it :)